OpenStudy (vortish):

Using the slope formula what is the slope of the line (2,-8) (8,-8) (a) the slope is blank (b) undifinded

OpenStudy (nubeer):

slope =(y2-y1)/(x2-x1)

OpenStudy (kira_yamato):

So continuing from where Itachi left off, slope =(-8-(-8))/(8-2) = 0/6 = 0 Slope = 0

OpenStudy (vortish):

can you expand on that a little more so is it undefinded or does it have a slope

OpenStudy (kira_yamato):

There is a slope 0